# A Token-level Text Image Foundation Model for Document Understanding

### v1

@@ Abstract
In recent years, general visual foundation models (VFMs) have witnessed increasing adoption, particularly as image encoders for popular multi-modal large language models (MLLMs). However, without semantically fine-grained supervision, these models still encounter fundamental prediction errors in the context of downstream text-image-related tasks, i.e., perception, understanding and reasoning with images containing small and dense texts. To bridge this gap, we develop TokenOCR, the first token-level visual foundation model specifically tailored for text-image-related tasks, designed to support a variety of traditional downstream applications. To facilitate the pretraining of TokenOCR, we also devise a high-quality data production pipeline that constructs the first token-level image text dataset, TokenIT, comprising 20 million images and 1.8 billion token-mask pairs. Furthermore, leveraging this foundation with exceptional image-as-text capability, we seamlessly replace previous VFMs with TokenOCR to construct a document-level MLLM, TokenVL, for VQA-based document understanding tasks. Finally, extensive experiments demonstrate the effectiveness of TokenOCR and TokenVL. Code, datasets, and weights will be available at https://token-family.github.io/TokenOCR_project/.
@@ 1 Introduction
TokenOCR token-level TokenIT 20M 1.8B
Figure 1: For different tasks, previous works select different VFMs from general foundation models (path 1). In contrast, we develop a unified token-level foundation model, TokenOCR, specifically tailored for text-image-related tasks (path 2). TokenOCR is trained on a substantial self-built dataset, TokenIT, comprising 20 million images and 1.8 billion token-mask pairs. This well-learned model is capable of supplanting other VFMs in related downstream tasks.
@@
Leveraging the self-constructed TokenIT dataset, we further propose the first token-level text image foundation model, named TokenOCR, designed to support a wide array of text-image-related downstream tasks. To achieve image-as-text semantic alignment, token-level visual embeddings are aligned with token-level language embeddings for positive token-mask pairs, meanwhile ensuring that negative pairs remain distinct within the embedding space. Specifically, each token-level visual embedding is derived through a mean pooling operation applied to the visual image features within a corresponding token mask; each token-level language embedding is produced via a straightforward token embedding layer, obviating the need for a complex text encoder like CLIP.
The image-as-text semantic attributes, aligned at the VFM level, effectively bridge the gaps between visual and language modalities. This approach creates a unified sequence representation that can be seamlessly integrated into any large language model (LLM) for popular MLLM tasks. Building upon this foundation, we propose a document-level MLLM, named TokenVL, which further enhances spatially visual-language token alignment at the LLM level for document understanding in Visual Question Answering (VQA) tasks. Additionally, we freeze the weights of the TokenOCR model to facilitate other downstream applications, including text segmentation, text retrieval, and end-to-end text recognition tasks.
@@
    The first token-level text image foundation model, TokenOCR, is proposed to support various downstream tasks, including text segmentation, text retrial, and text understanding.
@@
    Extensive experiments demonstrate the effectiveness of our proposed TokenOCR and TokenVL. Specifically, TokenOCR shows exceptional ”zero-shot” capabilities and flexibility compared to other VFMs, such as CLIP, SAM, and InternViT2.5. TokenVL with 8B parameters, incorporating TokenOCR as the VFM, achieves performance gains of 38 on the OCRBench task and an average of 8.8% across ten document VQA tasks. Similarly, TokenVL with 2B parameters results in performance gains of 17 on the OCRBench task and an average of 13.34% on the ten VQA tasks.
@@ 2 Related Work
Figure 3: An overview of the proposed TokenOCR, where the token-level image features and token-level language features are aligned within the same semantic space. This “image-as-text” alignment seamlessly facilitates user-interactive applications, including text segmentation, retrieval, and visual question answering.
@@ 4 Methodology
Utilizing the TokenIT dataset with 1.8B token-mask pairs, we construct the first token-level OCR foundation model (TokenOCR) by token-level image-as-text alignment. For VQA-based document understanding downstream tasks, we employ the well-learned foundation model to construct an MLLM (TokenVL), which includes the following stages: 1) LLM-guided Token Alignment; 2) Supervised Instruction Tuning. Besides, we also freeze the foundation model (unless otherwise stated) to conduct other text-related downstream tasks, including text segmentation, text retrieval, and text understanding.
4.1 TokenOCR
@@ 4.2 TokenVL
The image-as-text semantic attributes inherently bridge the gaps between visual and language modalities, creating a unified sequence representation that LLM can effectively understand. Inspired by this, we employ the TokenOCR as the visual foundation model and further develop an MLLM, named TokenVL, tailored for document understanding. Following the previous training paradigm (Chen et al., 2024d; Lv et al., 2023; Wei et al., 2025; Hu et al., 2024c), TokenVL also includes two stages: 1) LLM-guided Token Alignment Training for text parsing tasks and 2) Supervised Instruction Tuning for VQA tasks.
Specifically, adopting the widely-used multi-scale adaptive cropping strategy (Ye et al., 2023b), the input image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$ is initially divided into several non-overlapping sub-images $\{\mathbf{X}_{i}\in\mathbb{R}^{\iota\times\iota\times 3}|i\in\{1,2,...,N\}\}$. By default, $\iota$ is set to 448 and $N$ does not exceed 6. Additionally, the original image $\mathbf{X}$ is resized to a global image $\mathbf{X}_{g}$ with the same size to preserve the overall layout. Subsequently, our proposed TokenOCR processes these images $\mathcal{X}=\{\mathbf{X}_{g},\mathbf{X}_{1},...,\mathbf{X}_{N}\}$ to produce their corresponding visual embeddings, denoted as =$\mathcal{F}=\{\mathbf{\tilde{F}}_{i}\in\mathbb{R}^{\frac{4\times\iota}{p}\times\frac{4\times\iota}{p}\times D}|i\in\{g,1,2,...,N\}\}$.
After that, for each visual image features $\mathbf{\tilde{F}}_{i}$ (global image and sub-images), we apply a token abstractor $\xi:\mathbb{R}^{\frac{4\times\iota}{p}\times\frac{4\times\iota}{p}\times D}\xrightarrow{}\mathbb{R}^{\frac{\iota}{p\times\frac{s}{4}}\times\frac{\iota}{p\times\frac{s}{4}}\times D}$ to adaptively extract a meaningful visual embedding within each window of shape $s\times s$, where $s$ is set to 4 in our experiment. Specifically, in addition to the original dictionary of the tokenizer, we define a special token <text> to obtain a learnable token embedding $\mathbf{e}_{s}\in\mathbb{R}^{1\times 1\times D}$. Benefiting from the priors of the TokenOCR, the special token embedding can easily learn robust representations to identify the most suitable visual embeddings within each window. Concretely, for each sub-image and global image, we first re-organize the shape of its visual embeddings $\mathbf{\tilde{F}}_{i}$ from $\frac{4\times\iota}{p}\times\frac{4\times\iota}{p}\times D$ to $(\frac{\iota}{p\times\frac{s}{4}})^{2}\times D\times s^{2}$. $\xi(\cdot)$ is then implemented as follows:
@@ 5 Experiments
Implementation Details. To pre-train the TokenOCR foundation model, we employ the AdamW optimizer alongside a cosine learning rate schedule, with a base learning rate set at 5e-4. The model undergoes pre-training for two epochs on the TokenIT dataset. For TokenVL, we leverage the well-trained TokenOCR as the visual foundation model and InternLM (Cai et al., 2024) as the language model. Specifically, during the LLM-guided token alignment stage, InternLM remains frozen while we train the TokenOCR and newly introduced token abstractor. This stage involves training for one epoch on the TokenIT dataset, utilizing a base learning rate of 2e-4. In the subsequent supervised instruction tuning stage, all parameters are fully trainable, with a base learning rate of 1e-5. These experiments are executed on 64 H800 GPUs. Additional implementation details about other downstream experiments are provided within each respective subtask and Supplementary Material.
5.1 Effectiveness of TokenOCR
@@
Text Segmentation: 1) Zero-shot Segmentation: We compute the similarity between visual and language features to get the segmentation results. For CLIP, in line with prior work, we select “text” as the language prompt, which has been proven to be the most effective (Yu et al., 2023). In our method, we use a space “ ” as the language prompt and then apply a negation operation to derive the foreground similarity map. 2) Linear Probe: We keep the VFM frozen and train a linear layer to perform segmentation. Based on the results shown in Table 1, TokenOCR demonstrates significant average performance improvement across various text segmentation tasks. In the zero-shot setting, TokenOCR-1024px achieves the highest average score of 34.59%, significantly outperforming CLIP-L by 18.78%. In the linear probe setting, TokenOCR again leads with an average score of 48.77%, showing considerable improvements over SAM-H and InternViT2.5.
@@
Visual Question Answering: To further explore the representation learning capabilities of VFMs, we keep them frozen and fine-tune Vicuna-7B (Zheng et al., 2023) as the language model to conduct the text-related VQA tasks. All comparison methods employ the same configuration—training data, test benchmarks, learnable parameters, and optimizer—to ensure a fair evaluation. As seen in Table 2, TokenOCR achieves the highest scores on popular benchmarks, outperforming SAM-H, CLIP-L, and InternViT2.5 by 46.39%, 9.85%, and 0.96%, respectively.
@@
Text Retrieval: We select representative models, CLIP and InternViT2.5, to compare with our proposed TokenOCR on a Chinese dataset and an English dataset. Specifically, all VFMs are frozen. We calculate the similarity maps between the visual embeddings (extracted from the VFM) of all retrieval images and the language embeddings of all queries. For linear probe experiments, we use the same training data and train a simple linear classifier to score each similarity map, assigning a 1 if the similarity score is greater than 0.5, and a 0 otherwise. Finally, mean Average Precision (mAP) is employed to evaluate the performance of each VFM. The comparison results show that using only a few parameters, TokenOCR can perform well. Specifically, our proposed TokenOCR can achieve an average score of 63.62% on bilingual tasks. Additionally, there remains significant room for improvement through specific designs and components in the future.
@@ 6 Conclusion
In the paper, we take a step towards constructing a fine-grained visual foundation model, and propose a series of token-level product families: TokenIT, TokenOCR, and TokenVL. We also explore the potential and effectiveness of TokenOCR and TokenVL at a sufficiently large scale. While this approach demonstrates good and consistent performance gains on downstream tasks, there remains significant room for improvement through effective training strategies or additional designs. Therefore, we hope these products will serve as easily reproducible baselines for more downstream tasks in the future.
@@ Appendix B Interactive Demo
1) Our foundation model TokenOCR can distinguish text and background areas well. This means that when using the foundation model for downstream tasks, we can remove redundant background features at very low cost;
2) For complex, dense, and small texts, TokenOCR still precisely perceive, such as “picture”(Code), “f”(Code), “19”(Table), “P”(Table), etc. Additionally, our TokenOCR still can capture punctuation marks, such as commas, periods, double quotes, etc;
3) Our TokenOCR also supports handwritten texts, such as “STE”(Document) and “USA”(Document). We will provide a demo address for users to experience it.
@@ C.1 Data Source
Document DocVQA (Mathew et al., 2021), InfographicsVQA (Mathew et al., 2022), KIE, KleisterCharity (Stanisławek et al., 2021), PubTabNet (Zhong et al., 2020), RVL-CDIP (Harley et al., ), VisualMRC (Tanaka et al., 2021), Docmatix (Laurençon et al., 2024a), IIT-CDIP (Xu et al., 2019), publaynet (Zhong et al., 2019), Synthdog-en (Kim et al., 2022), DocGenome (Xia et al., 2024), CCpdf (Turski et al., 2023)
@@ C.2 Data Generation
2) Text Recognition. We use the previous SOTA method (Guan et al., 2023a) to obtain the recognition results for all types, except for natural scene text images. As these natural scene datasets already provide text transcriptions, we adopt them directly;
@@ Appendix E Mainstream Benchmark Results
TokenVL-2B TokenOCR InternLM2.5-1.8B 89.9 61.0 81.1 76.4 77.10
@@
TokenVL-8B TokenOCR InternLM2.5-7B 94.2 76.5 86.6 79.9 84.30

### v2

@@ Abstract
In recent years, general visual foundation models (VFMs) have witnessed increasing adoption, particularly as image encoders for popular multi-modal large language models (MLLMs). However, without semantically fine-grained supervision, these models still encounter fundamental prediction errors in the context of downstream text-image-related tasks, i.e., perception, understanding and reasoning with images containing small and dense texts. To bridge this gap, we develop TokenFD, the first token-level visual foundation model specifically tailored for text-image-related tasks, designed to support a variety of traditional downstream applications. To facilitate the pretraining of TokenFD, we also devise a high-quality data production pipeline that constructs the first token-level image text dataset, TokenIT, comprising 20 million images and 1.8 billion token-mask pairs. Furthermore, leveraging this foundation with exceptional image-as-text capability, we seamlessly replace previous VFMs with TokenFD to construct a document-level MLLM, TokenVL, for VQA-based document understanding tasks. Finally, extensive experiments demonstrate the effectiveness of TokenFD and TokenVL.
@@ 1 Introduction
TokenFD token-level TokenIT 20M 1.8B
Figure 1: For different tasks, previous works select different VFMs from general foundation models (path 1). In contrast, we develop a unified token-level foundation model, TokenFD, specifically tailored for text-image-related tasks (path 2). TokenFD is trained on a substantial self-built dataset, TokenIT, comprising 20 million images and 1.8 billion token-mask pairs. This well-learned model is capable of supplanting other VFMs in related downstream tasks.
@@
Leveraging the self-constructed TokenIT dataset, we further propose the first token-level text image foundation model, named TokenFD, designed to support a wide array of text-image-related downstream tasks. To achieve image-as-text semantic alignment, token-level visual embeddings are aligned with token-level language embeddings for positive token-mask pairs, meanwhile ensuring that negative pairs remain distinct within the embedding space. Specifically, each token-level visual embedding is derived through a mean pooling operation applied to the visual image features within a corresponding token mask; each token-level language embedding is produced via a straightforward token embedding layer, obviating the need for a complex text encoder like CLIP.
The image-as-text semantic attributes, aligned at the VFM level, effectively bridge the gaps between visual and language modalities. This approach creates a unified sequence representation that can be seamlessly integrated into any large language model (LLM) for popular MLLM tasks. Building upon this foundation, we propose a document-level MLLM, named TokenVL, which further enhances spatially visual-language token alignment at the LLM level for document understanding in Visual Question Answering (VQA) tasks. Additionally, we freeze the weights of the TokenFD model to facilitate other downstream applications, including text segmentation, text retrieval, and end-to-end text recognition tasks.
@@
    The first token-level text image foundation model, TokenFD, is proposed to support various downstream tasks, including text segmentation, text retrial, and text understanding.
@@
    Extensive experiments demonstrate the effectiveness of our proposed TokenFD and TokenVL. Specifically, TokenFD shows exceptional ”zero-shot” capabilities and flexibility compared to other VFMs, such as CLIP, SAM, and InternViT2.5. TokenVL with 8B parameters, incorporating TokenFD as the VFM, achieves performance gains of 38 on the OCRBench task and an average of 8.8% across ten document VQA tasks. Similarly, TokenVL with 2B parameters results in performance gains of 17 on the OCRBench task and an average of 13.34% on the ten VQA tasks.
@@ 2 Related Work
Figure 3: An overview of the proposed TokenFD, where the token-level image features and token-level language features are aligned within the same semantic space. This “image-as-text” alignment seamlessly facilitates user-interactive applications, including text segmentation, retrieval, and visual question answering.
@@ 4 Methodology
Utilizing the TokenIT dataset with 1.8B token-mask pairs, we construct the first token-level OCR foundation model (TokenFD) by token-level image-as-text alignment. For VQA-based document understanding downstream tasks, we employ the well-learned foundation model to construct an MLLM (TokenVL), which includes the following stages: 1) LLM-guided Token Alignment; 2) Supervised Instruction Tuning. Besides, we also freeze the foundation model (unless otherwise stated) to conduct other text-related downstream tasks, including text segmentation, text retrieval, and text understanding.
4.1 TokenFD
@@ 4.2 TokenVL
The image-as-text semantic attributes inherently bridge the gaps between visual and language modalities, creating a unified sequence representation that LLM can effectively understand. Inspired by this, we employ the TokenFD as the visual foundation model and further develop an MLLM, named TokenVL, tailored for document understanding. Following the previous training paradigm (Chen et al., 2024d; Lv et al., 2023; Wei et al., 2025; Hu et al., 2024c), TokenVL also includes two stages: 1) LLM-guided Token Alignment Training for text parsing tasks and 2) Supervised Instruction Tuning for VQA tasks.
Specifically, adopting the widely-used multi-scale adaptive cropping strategy (Ye et al., 2023b), the input image $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$ is initially divided into several non-overlapping sub-images $\{\mathbf{X}_{i}\in\mathbb{R}^{\iota\times\iota\times 3}|i\in\{1,2,...,N\}\}$. By default, $\iota$ is set to 448 and $N$ does not exceed 6. Additionally, the original image $\mathbf{X}$ is resized to a global image $\mathbf{X}_{g}$ with the same size to preserve the overall layout. Subsequently, our proposed TokenFD processes these images $\mathcal{X}=\{\mathbf{X}_{g},\mathbf{X}_{1},...,\mathbf{X}_{N}\}$ to produce their corresponding visual embeddings, denoted as =$\mathcal{F}=\{\mathbf{\tilde{F}}_{i}\in\mathbb{R}^{\frac{4\times\iota}{p}\times\frac{4\times\iota}{p}\times D}|i\in\{g,1,2,...,N\}\}$.
After that, for each visual image features $\mathbf{\tilde{F}}_{i}$ (global image and sub-images), we apply a token abstractor $\xi:\mathbb{R}^{\frac{4\times\iota}{p}\times\frac{4\times\iota}{p}\times D}\xrightarrow{}\mathbb{R}^{\frac{\iota}{p\times\frac{s}{4}}\times\frac{\iota}{p\times\frac{s}{4}}\times D}$ to adaptively extract a meaningful visual embedding within each window of shape $s\times s$, where $s$ is set to 4 in our experiment. Specifically, in addition to the original dictionary of the tokenizer, we define a special token <text> to obtain a learnable token embedding $\mathbf{e}_{s}\in\mathbb{R}^{1\times 1\times D}$. Benefiting from the priors of the TokenFD, the special token embedding can easily learn robust representations to identify the most suitable visual embeddings within each window. Concretely, for each sub-image and global image, we first re-organize the shape of its visual embeddings $\mathbf{\tilde{F}}_{i}$ from $\frac{4\times\iota}{p}\times\frac{4\times\iota}{p}\times D$ to $(\frac{\iota}{p\times\frac{s}{4}})^{2}\times D\times s^{2}$. $\xi(\cdot)$ is then implemented as follows:
@@ 5 Experiments
Implementation Details. To pre-train the TokenFD foundation model, we employ the AdamW optimizer alongside a cosine learning rate schedule, with a base learning rate set at 5e-4. The model undergoes pre-training for two epochs on the TokenIT dataset. For TokenVL, we leverage the well-trained TokenFD as the visual foundation model and InternLM (Cai et al., 2024) as the language model. Specifically, during the LLM-guided token alignment stage, InternLM remains frozen while we train the TokenFD and newly introduced token abstractor. This stage involves training for one epoch on the TokenIT dataset, utilizing a base learning rate of 2e-4. In the subsequent supervised instruction tuning stage, all parameters are fully trainable, with a base learning rate of 1e-5. These experiments are executed on 64 H800 GPUs. Additional implementation details about other downstream experiments are provided within each respective subtask and Supplementary Material.
5.1 Effectiveness of TokenFD
@@
Text Segmentation: 1) Zero-shot Segmentation: We compute the similarity between visual and language features to get the segmentation results. For CLIP, in line with prior work, we select “text” as the language prompt, which has been proven to be the most effective (Yu et al., 2023). In our method, we use a space “ ” as the language prompt and then apply a negation operation to derive the foreground similarity map. 2) Linear Probe: We keep the VFM frozen and train a linear layer to perform segmentation. Based on the results shown in Table 1, TokenFD demonstrates significant average performance improvement across various text segmentation tasks. In the zero-shot setting, TokenFD-1024px achieves the highest average score of 34.59%, significantly outperforming CLIP-L by 18.78%. In the linear probe setting, TokenFD again leads with an average score of 48.77%, showing considerable improvements over SAM-H and InternViT2.5.
@@
Visual Question Answering: To further explore the representation learning capabilities of VFMs, we keep them frozen and fine-tune Vicuna-7B (Zheng et al., 2023) as the language model to conduct the text-related VQA tasks. All comparison methods employ the same configuration—training data, test benchmarks, learnable parameters, and optimizer—to ensure a fair evaluation. As seen in Table 2, TokenFD achieves the highest scores on popular benchmarks, outperforming SAM-H, CLIP-L, and InternViT2.5 by 46.39%, 9.85%, and 0.96%, respectively.
@@
Text Retrieval: We select representative models, CLIP and InternViT2.5, to compare with our proposed TokenFD on a Chinese dataset and an English dataset. Specifically, all VFMs are frozen. We calculate the similarity maps between the visual embeddings (extracted from the VFM) of all retrieval images and the language embeddings of all queries. For linear probe experiments, we use the same training data and train a simple linear classifier to score each similarity map, assigning a 1 if the similarity score is greater than 0.5, and a 0 otherwise. Finally, mean Average Precision (mAP) is employed to evaluate the performance of each VFM. The comparison results show that using only a few parameters, TokenFD can perform well. Specifically, our proposed TokenFD can achieve an average score of 63.62% on bilingual tasks. Additionally, there remains significant room for improvement through specific designs and components in the future.
@@ 6 Conclusion
In the paper, we take a step towards constructing a fine-grained visual foundation model, and propose a series of token-level product families: TokenIT, TokenFD, and TokenVL. We also explore the potential and effectiveness of TokenFD and TokenVL at a sufficiently large scale. While this approach demonstrates good and consistent performance gains on downstream tasks, there remains significant room for improvement through effective training strategies or additional designs. Therefore, we hope these products will serve as easily reproducible baselines for more downstream tasks in the future.
@@ Appendix B Interactive Demo
1) Our foundation model TokenFD can distinguish text and background areas well. This means that when using the foundation model for downstream tasks, we can remove redundant background features at very low cost;
2) For complex, dense, and small texts, TokenFD still precisely perceive, such as “picture”(Code), “f”(Code), “19”(Table), “P”(Table), etc. Additionally, our TokenFD still can capture punctuation marks, such as commas, periods, double quotes, etc;
3) Our TokenFD also supports handwritten texts, such as “STE”(Document) and “USA”(Document). We will provide a demo address for users to experience it.
@@ C.1 Data Source
Document DocVQA (Mathew et al., 2021), InfographicsVQA (Mathew et al., 2022), KIE, KleisterCharity (Stanisławek et al., 2021), PubTabNet (Zhong et al., 2020), RVL-CDIP (Harley et al., 2015), VisualMRC (Tanaka et al., 2021), Docmatix (Laurençon et al., 2024a), IIT-CDIP (Xu et al., 2019), publaynet (Zhong et al., 2019), Synthdog-en (Kim et al., 2022), DocGenome (Xia et al., 2024), CCpdf (Turski et al., 2023)
@@ C.2 Data Generation
2) Text Recognition. We use the previous SOTA method (Guan et al., 2023b) to obtain the recognition results for all types, except for natural scene text images. As these natural scene datasets already provide text transcriptions, we adopt them directly;
@@ Appendix E Mainstream Benchmark Results
TokenVL-2B TokenFD InternLM2.5-1.8B 89.9 61.0 81.1 76.4 77.10
@@
TokenVL-8B TokenFD InternLM2.5-7B 94.2 76.5 86.6 79.9 84.30